Given that the data below was randomly selected from a lognormal distribution. What is the probability of a randomly selected item having a value greater than 3.9? Give your answer to 3 decimal places.

Data

1.72, 124.2, 1.04, 0.84, 12.82, 11.76, 51.2, 0.11
Answer #1

from above mean =25.46

sample variance =1881.184

therefore parameter for normal transformation:

std dev(ω)= √(ln((σ22)+1))= 1.1668
mean(ϴ)= ln(μ/(√((σ22)+1)))= 2.5564
probability =P(X>3.9)=1-P(X<3.9)=1-P(Z<(ln(3.9)-2.56)/1.1668)=1-P(Z<-1.02)=0.846